OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

so do you know the formula for equation of line if (x1,y1) = (5,-2) and m = 2/5 ?

OpenStudy (anonymous):

The formula is y-2=2/5(x-5)